# ProSGNeRF: Progressive Dynamic Neural Scene Graph with Frequency Modulated Auto-Encoder in Urban Scenes

### 14 Dec 2023 v1

@@ 1 Introduction
Figure 1: The isometric view of the proposed method, ProSGNeRF. We employed a 2D segmentation network, SAM, to preprocess the training data and generate accurate masks for dynamic objects. We propose a progressive neural scene graph approach that dynamically allocates local neural scene graph (box). The entire scene is decomposed into three parts: background, dynamic objects, and far-field. We design separate networks for background and objects and introduce a far-field loss for regularization. Nodes $l$ represent individual dynamic objects. $F_{bkg}$ models the static background scene and $F_{obj}$ models movable foreground objects in local object-centric coordinate frames.
@@
  • We propose a novel editable neural scene rendering method, ProSGNeRF, which decomposes dynamic, background, and far-field into a progressively neural scene graph with decoupled object transformations and scene representations and also supports object manipulation and scene roaming.
@@ 2 Related work
NSG [16] and SUDS [21] are the most relevant works to ours. NSG adopts neural scene graphs and dynamic object bounding boxes to decouple the scene but cannot handle ego-motion. Besides, NSG needs over 1TB of memory to represent a 30-second video. SUDS [21] uses a large model (DINO [4]) to distill scene features and obtain semantic information and scene flow. Their methods can handle camera ego-motion and static objects but struggle with high-dynamic objects. In contrast, our approach effectively addresses large-scale camera ego-motion and dynamic objects, enabling urban scene editing useful for constructing simulators for autonomous driving.
@@ 3 Method
The pipeline of our system is shown in Fig. 1. We use a set of sequential RGB images $\{I_{i}\}_{i=1}^{M}$ with known pose $\{R|t\}$ as our input. We construct the progressive neural scene graph with the instance mask from SAM and 3D bounding box from machine-predicted. The scene is decomposed into multi-local scene graphs with three branches: background, dynamic objects, and far-field. In the dynamic objects branch, we design a frequency auto-encoder to learn a normalized, object-centric representation whose embedding describes and disentangles shape, appearance, and pose. The decoder translates the features into occupancy and color information for given 3D points and viewing directions in object space. We also design lidar points projection geometry consistency in large-scale urban scenes.
Figure 2: The isometric view of the proposed progressive neural scene graph. This scene representation uses a progressive scheme that dynamically allocates a local neural scene graph (box) based on the camera pose. Adjacent local scene graphs have overlapping regions to maintain global consistency. The leaf nodes $l$ are visualized as boxes with their local Cartesian coordinate axis. We also visualize the transformations and scaling between root and leaf coordinate frames using arrows with annotated transformation and scale matrices. The overall representation model is denoted as $F_{\theta}$.
@@ 3.1 Progressive Neural Scene Graph
To this end, we design a progressive neural scene graph to represent city-scale scenes with multi-dynamic objects. The scene graph $\mathrm{SG}$, illustrated in Fig. 2, is composed of cameras, static nodes and a set of dynamic nodes which represent the dynamic components of the scene, including the object’s appearance, shape, and classes. The entire scene is decomposed into several local scene representations:
@@
For all edges originating from the root node $\mathcal{W}$, we assign transformations between the global world space and the local objects with $T$. We compute the scaling factor $S_{o}$, assigned at edge $e_{o}$ between an object node and its shared representation model $F$.
@@ 3.3 Frequency Auto-Encoder
Figure 3: Qualitative results on reconstruction and novel scene arrangements of a scene from the KITTI dataset [8] for NeRF [14], NSG [16], SUDS [21] and our method. From left to right, these images correspond to different timesteps captured in the dynamic scene. We place the PSNR values of each scene in the bottom right corner. NeRF and SUDS are limited in their ability to properly represent the dynamic parts of the scene. In contrast, our neural scene graph method achieves high-quality view synthesis results for both reconstruction and novel scene synthesis, regardless of scene dynamics.
@@ 4.1 Experimental Results
We then generate previously unseen frames of novel object arrangements, temporal variations. Our progressive neural scene graph architecture allows dynamic urban scene editing, which can be manipulated at its edges and nodes. To demonstrate the flexibility of the proposed scene representation, we rotate the pose of dynamic objects. We present the view synthesis results in Fig. 4. Note that we have not observed the rotation of dynamic objects during training.
@@ 4.2 Ablation Study
w/o progressive scene graph 27.21 0.782 0.131
w/o frequency auto-encoder 28.32 0.823 0.109
w/o depth loss 28.93 0.824 0.133
w/o KL divergence loss 29.45 0.84 0.128
Full Model 30.31 0.891 0.098

### 15 Dec 2023 v2

@@ 1 Introduction
Figure 1: The isometric view of the proposed method, ProSGNeRF. We employed a 2D segmentation network, SAM, to preprocess the training data and generate accurate masks for dynamic objects. We propose a progressive neural scene graph approach that dynamically allocates local scene graph (box). The entire scene is decomposed into three parts: background, dynamic objects, and far-field. We design separate networks for background and objects and introduce a far-field loss for regularization. Nodes $l$ represent individual dynamic objects. $F_{bkg}$ models the static background scene and $F_{obj}$ models movable foreground objects in local object-centric coordinate frames.
@@
  • We propose a novel editable scene reconstruction method, ProSGNeRF, which decomposes dynamic, background, and far-field into a progressively neural scene graph. Our method also supports object manipulation and scene roaming.
@@ 2 Related work
NSG [16], SUDS [21], and Mars [22] are the most relevant works to ours. Both NSG and Mars adopts neural scene graphs and dynamic object bounding boxes to decouple the scene but cannot handle ego-motion. Besides, NSG needs over 1TB of memory to represent a 30-second video. SUDS [21] uses a large model (DINO [4]) to distill scene features and obtain semantic information and scene flow. Their methods can handle camera ego-motion and static objects but struggle with high-dynamic objects. In contrast, our approach effectively addresses large-scale camera ego-motion and dynamic objects, enabling urban scene editing useful for constructing simulators for autonomous driving.
@@ 3 Method
The pipeline of our system is shown in Fig. 1. We use a set of sequential RGB images $\{I_{i}\}_{i=1}^{M}$ with known pose $\{R|t\}$ as our input. We construct the progressive neural scene graph with the instance mask from SAM and 3D bounding box. The scene is decomposed into multi-local scene graphs with three branches: background, dynamic objects, and far-field. In the dynamic objects branch, we design a frequency auto-encoder to learn a normalized, object-centric representation whose embedding describes and disentangles shape, appearance, and pose. The decoder translates the features into occupancy and color information for given 3D points and viewing directions in object space. We also design lidar points projection geometry consistency in large-scale urban scenes.
[…]
@@ 3.1 Progressive Neural Scene Graph
To this end, we design a progressive neural scene graph to represent city-scale scenes with multi-dynamic objects. In Figure 2, the scene graph $\mathrm{SG}$ is depicted, encompassing cameras, static nodes, and a collection of dynamic nodes. These dynamic nodes capture the evolving aspects of the scene, encapsulating the object’s appearance, shape, and various classes.
The entire scene is decomposed into several local scene representations:
@@
[…]
@@ 3.3 Frequency Auto-Encoder
Figure 3: Qualitative results on reconstruction and novel scene arrangements of a scene from the KITTI dataset [8] for NeRF [14], NSG [16], SUDS [21] and our method. From left to right, these images correspond to different timesteps captured in the dynamic scene. We place the PSNR values of each scene in the bottom right corner. NeRF and SUDS are limited in their ability to properly represent the dynamic parts of the scene. In contrast, our neural scene graph method achieves high-quality view synthesis results for both image reconstruction and novel scene synthesis, regardless of scene dynamics.
@@ 4.1 Experimental Results
We then generate previously unseen frames of novel object arrangements, temporal variations. Our progressive neural scene graph architecture allows dynamic urban scene editing, which can be manipulated at its edges and nodes. To exhibit the flexibility of our proposed scene representation, we manipulate the pose of dynamic objects through rotation. We present the view synthesis results in Fig. 4. Note that we have not observed the rotation of dynamic objects during training.
@@ 4.2 Ablation Study
w/o progressive scene graph 27.21 0.782 0.079
w/o frequency auto-encoder 28.32 0.823 0.069
w/o depth loss 28.93 0.824 0.078
w/o KL divergence loss 29.45 0.84 0.073
Full Model 30.31 0.891 0.055
@@ 5 Conclusion
6 Acknowledgement
Authors gratefully appreciate the contribution of Yanbo Wang from Shanghai Jiao Tong University, Zirui Wu and Yongliang Shi from Institute for AI Industry Research (AIR), Tsinghua University.